OpenStudy (anonymous):

1. A sample of gold (Au) has a mass of 35.12 g. a. Calculate the number of moles of gold (Au) in the sample. . b. Calculate the number of atoms of gold (Au) in the sample

OpenStudy (michele_laino):

please in order to calculate number of moles, you have to divide your mass, by atomic mass of Au, please keep in mind that atomic mass of Au =197, so?

OpenStudy (michele_laino):

moles of Au=\[\frac{ 35.12 }{ 197 }=...\] please continue @denii_114

OpenStudy (anonymous):

Would it be 0.178??

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

please try, it is a constant whose value is very huge

OpenStudy (michele_laino):

I lost the connection, I repeat: how many atoms of Au there are in a mole of Au? that number is very huge, its name is Avogadro constant

OpenStudy (anonymous):

Oh.... Um 79?

OpenStudy (michele_laino):

pratically to answer to your second question, you have to multiply numbers of your moles of Au by the Avogadro constant. Please do you know what is the value of Avogadro constant?

OpenStudy (anonymous):

No, not really...

OpenStudy (michele_laino):

Avogadro constant is: \[N _{A}=6.02*10^{23}\]

OpenStudy (michele_laino):

so yo\[N _{A}*0.178=6.023*10^{23}*0.178=...\]u have to calculate this: and you will get your number of atoms of Au, please try!

OpenStudy (anonymous):

Okay, Please excuse me if im wrong, im tryin.... Wld it be... 1.072??

OpenStudy (anonymous):

I was absent the day the we're showing it... Doing the best i can.

OpenStudy (michele_laino):

your answer is : \[1.072*10^{23}\] atoms of Au

OpenStudy (anonymous):

Oh Okay.... So from the answer i get, i always add in the \[10^{23}\]?

OpenStudy (michele_laino):

yes! you have to multiply your answer 1.072 by the exponential factor, don't forget it, please!

OpenStudy (anonymous):

Ohhh.. Okay, got it. Thank You so much!! @Michele_Laino

OpenStudy (michele_laino):

Thank you!

OpenStudy (anonymous):

Okay with that i cld answer the rest of the question similar to it. You think you can heklp me with one more thing, please?? @Michele_Laino

OpenStudy (michele_laino):

@denii_114 yes!

OpenStudy (anonymous):

Okay Thank You!! Here it is.... 2. A sample of table sugar (sucrose, C12H22O11) has a mass of 1.202 g. a. Calculate the number of moles of C12H22O11 contained in the sample b. Calculate the moles of each element in C12H22O11. Show your work. c. Calculate the number of atoms of each type in C12H22O11. Show your work. @Michele_Laino

OpenStudy (michele_laino):

please you have to calculate the mass of sucrose molecule. now a sucrose molecule has 12 aoms of C, 22 atoms of H, and 11 atoms of O

OpenStudy (anonymous):

Got it!

OpenStudy (michele_laino):

so we have to consider tha atomic mass of C, H and O, please take your periodic table and write the atomic mass of C, H and O

OpenStudy (michele_laino):

from mine, I read: atomic mass of C=12 please continue

OpenStudy (anonymous):

H=22 O=11 ?

OpenStudy (michele_laino):

please you have to read atomic masses from periodic table of elements, do you have it, on your chemistry textbook?

OpenStudy (anonymous):

I'll look online for one...

OpenStudy (michele_laino):

ok! atomic mass of H=1.008, which can be rounded to 1

OpenStudy (anonymous):

H=1 o=15

OpenStudy (michele_laino):

please you have to round off, we have O=15.99, namely 16. are you agree?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Forgot to round sorry

OpenStudy (michele_laino):

ok! Now in order to find the mass of sucrose molecule, you have to do this calculus: \[12*2+22*1+16*11\]

OpenStudy (michele_laino):

sorry I have made an error, I rewrite: \[12*12+22*1+11*16=...\]

OpenStudy (anonymous):

222, right?

OpenStudy (anonymous):

Wait. did sumthing wrong

OpenStudy (anonymous):

342

OpenStudy (michele_laino):

perfect!

OpenStudy (michele_laino):

in order to find the sucrose moles, you have to do like previous exercise, namely you have to divide the mass of sucrose, by the molecule mass of sucrose: \[\frac{ 1.202 }{ 342 }=...\]

OpenStudy (michele_laino):

please try?

OpenStudy (anonymous):

do i round too??

OpenStudy (anonymous):

i am trying...

OpenStudy (anonymous):

Im doing it... this stupid keyboard takes forever to write in..... I got 0.003.... round? 0.004.

OpenStudy (michele_laino):

are you sure? please retry

OpenStudy (michele_laino):

sorry I have made an error, you are right!

OpenStudy (anonymous):

yeaa...

OpenStudy (michele_laino):

ok! now you know that one sucrose mole contains 6.02*10^23 sucrose molecules, but we have only 0.004 molecules , namely 4*10^-3, so we have to multiply 0.004 by Avogadro constant, like the previous exercise, so: \[0.004*6.02*10^{23}=...\]

OpenStudy (anonymous):

So how do you find the moles for each element?? Same as first question?

OpenStudy (anonymous):

Oh okay....

OpenStudy (anonymous):

2.41

OpenStudy (michele_laino):

I got 0.0241*10^23

OpenStudy (anonymous):

one question... you get the "0.0241"? if it was 2.41

OpenStudy (anonymous):

**how you...

OpenStudy (michele_laino):

why? you have to find 0.004*6.023*10^23

OpenStudy (anonymous):

I keep getting 2.40

OpenStudy (michele_laino):

sorry, please find 0.004*6.023=...

OpenStudy (anonymous):

Okay.. NOW i get 0.024...

OpenStudy (michele_laino):

ok! now you have to multiply 0.024 by 10^23, so you should get: \[0.024*10^{23}=2.4*10^{-2}*10^{23}=2.4*10^{-2+23}=2.4*10^{21}\]

OpenStudy (anonymous):

Yeaa thts what i did.. i was one step ahead of you...i got tht the first time

OpenStudy (michele_laino):

ok now we know that we have 2.4*10^21 molecules of sucrose. Please note that in one molecule of sucrose we have 12 atoms of C, so how many atoms of C we have in 2.4 molecules of sucrose?

OpenStudy (michele_laino):

...sorry in 2.4*10^21 molecules of sucrose...

OpenStudy (anonymous):

Do i add them into 2.4 * 10^21?

OpenStudy (michele_laino):

you have to multiply 2.4*10^21 by 12

OpenStudy (anonymous):

What i meant. Okay i got... 2.88

OpenStudy (michele_laino):

please find 2.4*12...

OpenStudy (anonymous):

28.8

OpenStudy (michele_laino):

don't forget the exponential factor....

OpenStudy (anonymous):

2.88

OpenStudy (anonymous):

Do i do tht for the rest?? H & O?? H= 2.4 O= 3.6

OpenStudy (michele_laino):

please, note that we have 28.8 *10^21 atoms of C, 52.8*10^21 atoms of H, 26.4*10^21 atoms of O are you agree?

OpenStudy (anonymous):

Yeah...

OpenStudy (michele_laino):

that answers to c.

OpenStudy (anonymous):

Yup got it...

OpenStudy (michele_laino):

we have to answer to b.

OpenStudy (michele_laino):

now, please write the value of the mass of an atom of C, please

OpenStudy (anonymous):

12

OpenStudy (michele_laino):

you are right! but I want to know how many grams!

OpenStudy (anonymous):

.100g

OpenStudy (michele_laino):

It is very simple, you have to multiply 12 by another constant, whose value is 1.66*10^-24, so we have: \[12*1.66*10^{-24}\]

OpenStudy (michele_laino):

is it clear?

OpenStudy (anonymous):

No, where you get the 1.66??

OpenStudy (michele_laino):

it is another constant whose name is "Dalton", it is tabulated in all chemical tables. In general in ordr to find a mass (in grams) of an atom, you have to multiply tha atomic mass of that element by the "Dalton"

OpenStudy (anonymous):

oh oh oh..... okay...

OpenStudy (michele_laino):

for example, atomic mass of Fe is 58.6, then the mass of atoms of Fe, expressed in grams is: \[58.6*1.66*10^{-24}=97.3*10^-24 grams\]

OpenStudy (anonymous):

Oh okay.. so it be 19.92 * 10-24grams??

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

So tht would be for C... H would be .... 1.66 * 10-24??

OpenStudy (michele_laino):

now you have to find the mass (in grams) of 28.8*10^21 atoms of C, namely you have to multiply your number by 28.8*10^21

OpenStudy (michele_laino):

so?

OpenStudy (anonymous):

Multiply 19.92 * 28.8 ??

OpenStudy (michele_laino):

\[19.92*10^{-24}*28.8*10^{21}=....\]

OpenStudy (anonymous):

0.57??

OpenStudy (michele_laino):

that's right! we have 0.57 grams of C

OpenStudy (michele_laino):

and how many moles of C?

OpenStudy (anonymous):

0.47?